OpenStudy (anonymous):

If one side of a rectangle increases ar a rate of 2 m/s and the other side decreases at a rate of 3 m/s, find the rate of change of the area when the first side is 4 m and the second side is 5 m.

hero (hero):

Now you're doing calculus?

hero (hero):

Oh, nevermind, lol. You were always doing calculus

hero (hero):

A = xy A_t' = (dx/dt)y + x(dy/dt)

OpenStudy (anonymous):

area : oneside * otherside darea = dx*y + dy*x darea = 2*4 + 3*5 = 23

hero (hero):

A_t' = (2)4 + 3(5) A_t' = 8 + 15 = 23m^2

hero (hero):

23m/s^2

hero (hero):

Only thing cool sector....you're taking the derivative with respect to time

OpenStudy (anonymous):

so 23 is the answer? thats all there is to it?

OpenStudy (anonymous):

i was trying to make it way more complicated.